OpenStudy (lovelyanna):

Math problem for fun! :D I came across a math problem that I've never seen or done before, and was so curios to know how to solve it! Can someone please lend a hand? :) Problem below:

OpenStudy (lovelyanna):

@jigglypuff314

Atsie (atsie):

If you added 60 and 25 and 15 minutes up, wouldn't that be your answer? Or am I missing something?

jigglypuff314 (jigglypuff314):

http://purplemath.com/modules/workprob.htm

jigglypuff314 (jigglypuff314):

anna's got the right idea ^_^

jigglypuff314 (jigglypuff314):

except the numbers should be in relation to each other like how 60 is 4 times 15

Atsie (atsie):

Darn! I hate it when I thought I was close, and I end up not being so! xD The link you just gave is making my head spin now. :P

OpenStudy (lovelyanna):

I tried making the denominators the same for each fraction, but 25 doesn't work.

OpenStudy (lovelyanna):

Is there a different equation I can use?

jigglypuff314 (jigglypuff314):

wait! You were completely right about the 1/60 + 1/25 + 1/15 = 1/t that's all you needed ^_^

OpenStudy (lovelyanna):

Ok, that's good to know! But how do I find the answer?

OpenStudy (lovelyanna):

I need to add all of them up, but I can't do that without the denominators being the same.

jigglypuff314 (jigglypuff314):

You could try multiplying every term by 60t

OpenStudy (lovelyanna):

How do I do that...? :/

jigglypuff314 (jigglypuff314):

60t / 60 + 60t / 25 + 60t / 15 = 60t / t

OpenStudy (lovelyanna):

60t/60 = t, and 60t/25 = 12t/5. Wouldn't that make it even more complicated?

jigglypuff314 (jigglypuff314):

I think it would make it more managable? you can multiply again by 5

jigglypuff314 (jigglypuff314):

if you want to get rid of that one denominator

OpenStudy (lovelyanna):

Let me try. :)

OpenStudy (lovelyanna):

After simplifying and multiplying by 5, I got 5t + 12t + 25t = 300. Is that right?

jigglypuff314 (jigglypuff314):

60/15 = 4 right? then times 5

OpenStudy (lovelyanna):

I think I found an easier way. 300 can be the denominator for each fraction, right?

jigglypuff314 (jigglypuff314):

that's also right :)

OpenStudy (lovelyanna):

Awesome! I just got the answer :D

OpenStudy (lovelyanna):

Thanks for your help! :)

jigglypuff314 (jigglypuff314):

I'm glad I was able to help ^_^ (I was never really good at these work type problems xD)

OpenStudy (lovelyanna):

haha, you think I'm good at them? :P

jigglypuff314 (jigglypuff314):

yeah! You've gotten much better since the last time I helped you ;P

OpenStudy (lovelyanna):

ikr! Thanks! :D